A car travels 120 miles from A to B at 30 miles per hour but returns the same distance at 40 miles per hour. What is the average speed for the round trip?
Answer by richwmiller(17219)

first leg takes 4 hours and return takes 3 hours
240=r*7
240/7=r
34 2/7 mph average
not (40+30)/2=35 but is the extra work involved worth it?